OpenStudy (anonymous):

The function for the cost of materials to make a shirt is f(x) = five sixths x + 5, where x is the number of shirts. The function for the selling price of those shirts is g(f(x)), where g(x) = 5x + 6. Find the selling price of 18 shirts

OpenStudy (zzr0ck3r):

\(g(\color{blue}{f(x)})=g(\color{blue}{x+5})=5(\color{blue}{x+5})+6=5x+31\)

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (zzr0ck3r):

we need to now find \(g(f(18))=5*18+31\)

OpenStudy (anonymous):

121 @zzr0ck3r

OpenStudy (triciaal):

@zzr0ck3r f(x) =5/6 x+5

OpenStudy (anonymous):

these are the options 106 120 185 196 @triciall @zzr0ck3r

OpenStudy (triciaal):

@brangetsbuckets the method @zzr0ck3r used is fine but please verify the original f(x) I prefer to compute separate do f(x) then use this value to compute in g